# A Quadratically-Constrained Convex Approximation for the AC Optimal Power Flow

### v1

@@ Abstract
We introduce a quadratically-constrained approximation (QCAC) of the AC optimal power flow (AC-OPF) problem. Unlike existing approximations like the DC-OPF, our model does not rely on typical assumptions such as high reactance-to-resistance ratio, near-nominal voltage magnitudes, or small angle differences, and preserves the structural sparsity of the original AC power flow equations, making it suitable for decentralized power systems optimization problems. To achieve this, we reformulate the AC-OPF problem as a quadratically constrained quadratic program. The nonconvex terms are expressed as differences of convex functions, which are then convexified around a base point derived from a warm start of the nodal voltages. If this linearization results in a non-empty constraint set, the convexified constraints form an inner convex approximation. Our experimental results, based on Power Grid Library instances of up to 30,000 buses, demonstrate the effectiveness of the QCAC approximation with respect to other well-documented conic relaxations and a linear approximation. We further showcase its potential advantages over the well-documented second-order conic relaxation of the power flow equations in two proof-of-concept case studies: optimal reactive power dispatch in transmission networks and PV hosting capacity in distribution grids.
Index Terms:
AC optimal power flow, Convex approximation, Difference-of-convex programming, PV hosting capacity, Optimal reactive power dispatch, Taylor series.
Nomenclature
Sets:
@@
$\mathcal{E}$
Set of branches.
$\mathcal{E}_{i}$
Set of branches connected to node $i$.
Constants:
@@
$\overline{N}_{i}^{\rm sh}$
Maximum step of shunt compensation at node $i$.
@@
$\overline{P}_{k}/\underline{P}_{k}$
Maximum/Minimum active power output of unit $k$.
$\overline{Q}_{k}/\underline{Q}_{k}$
Maximum/Minimum reactive power output of unit $k$.
$\overline{S}_{ij}$
Maximum apparent power flow through the branch between nodes $i$ and $j$.
@@
$\overline{V}_{i}/\underline{V}_{i}$
Maximum/Minimum voltage magnitude at node $i$.
Variables:
@@
I Introduction
Power systems are undergoing important changes due to the pervasive electrification of transportation and industry sectors and the increasing share of renewable energy resources. System operators must rely on reliably solving the AC Optimal Power Flow (AC-OPF) problem to ensure economically efficient and secure planning and operation decisions. The main challenge of reliably solving the AC-OPF comes from the nonconvexity of the power flow equations, which describe the steady-state governing physical laws of the transmission network [1].
In recent years, significant efforts have been devoted to developing more tractable linear and convex relaxations of the power flow equations, which can be reliably solved using state-of-the-art optimization solvers. In particular, these efforts include the semi-definite programming [2], second-order cone [3], and the quadratic convex [4] relaxations. The key features of convex relaxations are their ability to provide infeasibility certificates, lower bounds on the optimal objective values, and globally optimal decisions when tight. The main drawbacks of using relaxations include the overestimation of power losses [5], relaxations of Ohm’s law [6], and their inability to leverage AC base solutions, also known as warm starts, from previously solved instances.
Another line of research is focused on developing tractable model-based approximations of the power flow equations [7]. One of the main advantages of such approximations is that the can (i) use solutions of previous operating conditions [8], (ii) leverage variable estimates from the solution of the state estimator for real-time applications, (iii) employ engineering assumptions [9], e.g., reactance-to-resistance ratio, flat voltage magnitudes, or angle differences, and/or (iv) reduce the approximation error based on iterative procedures [10]. More recently, these approximations can be applied to exploit the accuracy and inference speed of learning-based approaches, using the solution of historical (or synthetically generated) instances to predict base points [11].
Various approximations of the power flow equations rely on a truncated first-order Taylor series expansion of the nonconvex terms arising from the product of complex voltages. The difference between linear approximations using Taylor series lies in the selection of the independent variables, voltage magnitudes $v$ and phase angles $\theta$, to perform the truncated expansion. Zhang et al. [12] derives a linear power flow model using the $v$ and $\theta$ as independent variables. Yang et al. [13] refines such linear model by using $v^{2}$ and $\theta$, and consider the impact of $v$ on network losses in [14]. Fatemi et al. [15] use $v^{2}$ and $v^{2}\theta$ as independent variables and introduce an adjustment factor to minimize estimation errors. Li et al. in [16] propose a logarithmic transform of voltage magnitudes, preserving the impact of voltage magnitudes and network losses on the active and reactive load flows. Yang et al. [17] derive a general formulation, where the power injections are expressed as functions of the independent variables, for selecting independent variables that render a linear power flow with reduced errors. Fan et al. [18] formulate a model to find a linear model with reduced linearization error by selecting the optimal independent variables for each bus. Long et al. in [19] propose a clustering-based approach to select the best linear power flow model depending on the similarity of the linearization error features of different operating conditions.
The authors in [20] and [10] propose the IV-Flow model, a linear approximation of the power flow equations by formulating the active and reactive power flow equations in rectangular coordinates in terms of voltages and currents and linearizing them using a first-order Taylor series expansion. Such approximation, along with the linearization of other constraints of the AC-OPF problem, is iteratively solved until a convergence criterion is satisfied. Mhanna and Mancarella [21] propose a sequential linear programming approach based on the construction of supporting hyperplanes and halfspaces and the exclusive solution of linear programming problems. Coffrin and Van Hentenryck [8] propose the LPAC model, which relies on a convex approximation of the cosine terms in the power flow equations and a Taylor approximation of the remaining nonlinear terms. Bolognani and Dörfler [22] propose an implicit linearization of the power flow manifold with a bound on the worst-case error.
On the other hand, first-order Taylor series expansions can also be applied to derive convex approximations, where all the convexities are preserved in the model while the nonconvexities are linearized. Jabr in [3] proposes a second-order conic (SOC) approximation based on his SOC relaxation, where the quadratic equality constraints are relaxed into a convex inequality and the nonconvex (arctangent) constraints are linearized. Coffrin et al. in [23] derive a quadratic programming approximation based on a second-order Taylor series expansion of the sine and cosine functions, assuming a near-nominal voltage magnitude approximation and a small angle difference, and on the relaxation of the quadratic equality constraints into inequalities. Baradar and Hesamzadeh in [24] propose a conic approximation of the OPF problem, assuming near nominal voltage magnitudes and small angle differences, expressed in terms of voltage magnitudes, power injections, and power losses. Šepetanc and Pandžić in [25] propose a second-order Taylor approximation of the AC power flows in polar coordinates by distributing power losses to both branch ends, and they assess its accuracy in a unit commitment problem.
Merkli et al. [26] formulate the AC power flow problem as a nonconvex quadratically constrained linear problem aiming at minimizing the deviation of power generation from given setpoints. The problem is reformulated as a difference-of-convex functions programming problem and solved using a difference-of-convex functions algorithm. The main limitations of such work are threefold: (i) the computational complexity of splitting the indefinite matrices in the nonconvex quadratic constraints, which relies on their full eigendecomposition (which takes $\mathcal{O}(n^{3})$ time), (ii) the matrices to be split depend on the network topology, which makes such decomposition impractical in scenarios where the topology varies as it requires a new decomposition each time the topology is altered, and (iii) the convexified constraints do not preserve the sparsity of the original ones and are generally dense, which is computationally challenging for optimization solvers in terms of internal operations and memory usage.
To bridge this gap, this paper proposes a quadratically-constrained convex approximation of the AC-OPF problem (QCAC) that does not rely on typical engineering assumptions such as high reactance-to-resistance ratio, near-nominal voltage magnitudes, or small angle differences while preserving the structural sparsity of the original nonconvex constraints. The proposed QCAC approximation is derived by lifting the AC-OPF in rectangular coordinates, which allows a reformulating as a difference-of-convex functions programming problem. The reformulated problem problem is then convexified around a given set of nodal voltages. The proposed QCAC approximation shows two important features: (i) it retains all the convex information of the nonconvex voltage relationships and solely linearizes their concave term, unlike other linear approximations, and (ii) it does not make any additional assumptions in terms of angle differences, reactance-to-resistance ratio, or voltage magnitudes. The accuracy and computational performance of the proposed approximation is numerically tested using realistic test systems up to 30,000 buses and two proof-of-concept case studies in transmission and distribution grids.
The remainder of this paper is organized as follows: Section II presents the formulation of the AC-OPF problem in rectangular coordinates and derives the proposed QCAC approximation. Section III illustrates the numerical performance of the proposed approximation with respect to the SOC and SDP outer approximations and a first-order Taylor series linearization. Section IV demonstrates two case studies of potential applications of the proposed model. Section V concludes the paper.
II Proposed Quadratically-Constrained Convex Approximation of the AC-OPF Problem
Next, we present the general formulation of the AC-OPF problem in lifted rectangular coordinates and derive the proposed quadratically-constrained approximation.
II-A Problem Formulation
@@
II-B The QCAC Approximation of the AC Optimal Power Flow
@@
Note that Problem (3) is an exact reformulation of the AC-OPF problem, which is still a nonconvex problem due to the convex terms in the right-hand side of inequalities (3b)-(3f). The above reformulation admits various (piecewise-)linearizations since the nonconvexities in the right-hand side of the inequalities can be expressed as separable functions. The drawback of piecewise linearizations is the need to enforce SOS2 constraints, which renders a mixed-integer second-order conic problem. In this paper, to retain the continuous nature of the AC-OPF problem, we propose to linearize such terms using a first-order Taylor series approximation around given nodal voltages, $V_{i}^{\rm re}$ and $V_{i}^{\rm im}$, rendering the following QCAC approximation:
@@
$\;c_{ii}\leq 2\left(V_{i}^{\rm re}v_{i}^{\rm re}+V_{i}^{\rm im}v_{i}^{\rm im}\right)-$
$\;\quad\left(\Big{(}V_{i}^{\rm re}\Big{)}^{2}+\Big{(}V_{i}^{\rm im}\Big{)}^{2}\right)+\xi_{i}^{\rm c},\;\forall i\in\mathcal{N},$ (4c)
@@
II-C Advantages of the QCAC approximation
@@
    The convexified constraints retain more information with respect to other linear approximations since they preserve all the information of the convex component of each nonconvex term and only linearize its concave part [27]. Conversely, in the linearization-based approaches, the second-order information is the objective constraints and/or constraints is lost. Figure 1 illustrates in a simple difference-of-convex (nonconvex) constraint, $y\geq x^{2}-\lvert x\rvert$, the approximate convex set obtained using the proposed convex approximation and the one obtained a standard first-order Taylor series linearization.
@@
    The QCAC approximation can be used along learning-based techniques that can leverage historical (or synthetically generated) instances to learn the mapping between nodal demands and nodal voltages, which correspond to our convexification points [11]. Notice that since the QCAC approximation is convex, it can be embedded within the training pipeline of a learning model as a differentiable layer, enabling the use of packages like DiffOpt.jl [28] and CVXPYLayers [29] to automate the computation of the gradients.
@@
III Numerical experiments
In this section, we provide numerical experiments to evaluate the performance of the proposed QCAC approximation in terms of feasibility, optimality gap, and runtime. The data of the studied instances can be retrieved from the PGLib-OPF library [30]. The numerical experiments have been implemented on a Macbook Pro laptop with an Apple M2 Pro processor and 32 GB of RAM under JuMP 1.10 with Mosek 10.1.2, Ipopt, and Gurobi 11.0.2 as the optimization solvers for the convex, nonconvex, and mixed-integer nonconvex optimization problems, respectively.
III-A Experimental setup
The results of the proposed QCAC approximation are compared with those obtained using a first-order Taylor series linearization (TS) of constraints (1l)-(1n), as well as with a second-order conic (SOC) and a semidefinite (SDP) relaxation [3, 2], for selected systems with up to 1354 buses. For the QCAC and TS models, the linearization/convexification points are set to the solution of the base case. For larger systems of 6,495, 13,659, and 30,000 buses, we study the performance of the QCAC model with respect to the solution of the original problem. The nonconvex, SOC, and SDP models are solved in PowerModels.jl in JuMP [31]. For each system, we generate 100 demand samples by multiplying the base case’s active and reactive power demand by a random vector drawn from a normal distribution with mean $1$ and standard deviation $0.1$.
TABLE I: Optimality Gaps, Distances to Feasibility, and Runtimes for Selected PGLIB Test Cases
Optimality Gap (%) Distance to feasibility (p.u.) Runtime (s)
Test case QCAC TS SOC SDP QCAC TS SOC SDP QCAC TS SOC SDP
pglib_opf_case30 0.06256 0.27451 0.66478 0.00680 0.14951 0.39028 0.29270 0.00004 0.03 0.03 0.04 0.82
pglib_opf_case39 0.27498 2.60725 3.26863 0.21014 0.01517 0.80445 1.09835 0.07128 0.03 0.03 0.04 2.82
pglib_opf_case57 0.01706 0.05375 0.00402 0.00092 0.01557 0.61258 0.01900 0.01852 0.05 0.04 0.04 12.39
pglib_opf_case89 1.49676 7.79117 1.14178 7.15318 0.36341 1.29853 0.52210 1.18675 0.37 0.35 0.51 155.7
pglib_opf_case118 0.37494 10.01670 2.89311 2.06640 0.00300 0.33364 0.09849 0.06768 0.18 0.17 0.24 638.3
pglib_opf_case179 1.21317 0.04374 2.46625 - 0.02196 0.004736 1.32447 - 0.58 0.36 0.19 -
pglib_opf_case300 1.53902 6.82574 1.23162 - 0.18725 0.39919 0.48233 - 1.83 2.21 2.16 -
pglib_opf_case500 0.31590 2.06582 0.13161 - 0.01643 0.14463 0.01076 - 1.05 1.00 1.39 -
pglib_opf_case793 0.81594 21.87617 3.85507 - 0.00300 0.16165 0.09789 - 4.81 4.01 4.32 -
pglib_opf_case1354 1.35456 57.87368 4.12804 - 0.01238 1.61108 0.54144 - 4.32 4.32 5.09 -
The highlighted values correspond to the best out of the QCAC, TS, SOC, and SDP models.
III-B Metrics
III-B1 Optimality gap

### v2

@@ Abstract
We introduce a quadratically-constrained approximation (QCAC) of the AC optimal power flow (AC-OPF) problem. Unlike existing approximations like the DC-OPF, our model does not rely on typical assumptions such as high reactance-to-resistance ratio, near-nominal voltage magnitudes, or small angle differences, and preserves the structural sparsity of the original AC power flow equations, making it suitable for decentralized power systems optimization problems. To achieve this, we reformulate the AC-OPF problem as a quadratically constrained quadratic program. The nonconvex terms are expressed as differences of convex functions, which are then convexified around a base point derived from a warm start of the nodal voltages. If this linearization results in a non-empty constraint set, the convexified constraints form an inner convex approximation. Our experimental results, based on Power Grid Library instances of up to 30,000 buses, demonstrate the effectiveness of the QCAC approximation with respect to other well-documented conic relaxations and a Taylor-series linear approximation. We further showcase its potential advantages over the well-documented second-order conic relaxation of the power flow equations in the optimal reactive power dispatch and photovoltaic hosting capacity problem.
1 Introduction
Power systems are undergoing important changes due to the pervasive electrification of transportation and industry sectors and the increasing share of renewable energy resources. System operators must rely on reliably solving the AC Optimal Power Flow (AC-OPF) problem to ensure economically efficient and secure planning and operation decisions. The main challenge of reliably solving the AC-OPF comes from the nonconvexity of the power flow equations, which describe the steady-state governing physical laws of the transmission network Capitanescu2016 .
In recent years, significant efforts have been devoted to developing more tractable linear and convex relaxations of the power flow equations, which can be reliably solved using state-of-the-art optimization solvers. In particular, these efforts include the semi-definite programming Bai2008 , second-order cone Jabr2006 , and the quadratic convex Coffrin2016 relaxations. The key features of convex relaxations are their ability to provide infeasibility certificates, lower bounds on the optimal objective values, and globally optimal decisions when tight. The main drawbacks of using relaxations include the overestimation of power losses Chen2016 , relaxations of Ohm’s law Sojoudi2012 , and their inability to leverage AC base solutions, also known as warm starts, from previously solved instances.
Another line of research is focused on developing tractable model-based approximations of the power flow equations Molzahn2019 . One of the main advantages of such approximations is that the can (i) use solutions of previous operating conditions Coffrin2014 , (ii) leverage variable estimates from the solution of the state estimator for real-time applications, (iii) employ engineering assumptions Alsac1974 , e.g., reactance-to-resistance ratio, flat voltage magnitudes, or angle differences, and/or (iv) reduce the approximation error based on iterative procedures Castillo2016 . More recently, these approximations can be applied to exploit the accuracy and inference speed of learning-based approaches, using the solution of historical (or synthetically generated) instances to predict base points Pineda2024 .
Various approximations of the power flow equations rely on a truncated first-order Taylor series expansion of the nonconvex terms arising from the product of complex voltages. The difference between linear approximations using Taylor series lies in the selection of the independent variables, voltage magnitudes $v$ and phase angles $\theta$, to perform the truncated expansion. Zhang et al. Zhang2013 derives a linear power flow model using the $v$ and $\theta$ as independent variables. Yang et al. Yang2017b refines such linear model by using $v^{2}$ and $\theta$, and consider the impact of $v$ on network losses in Yang2018 . Fatemi et al. Fatemi2015 use $v^{2}$ and $v^{2}\theta$ as independent variables and introduce an adjustment factor to minimize estimation errors. Li et al. in Li2018 propose a logarithmic transform of voltage magnitudes, preserving the impact of voltage magnitudes and network losses on the active and reactive load flows. Yang et al. Yang2019 derive a general formulation, where the power injections are expressed as functions of the independent variables, for selecting independent variables that render a linear power flow with reduced errors. Fan et al. Fan2021 formulate a model to find a linear model with reduced linearization error by selecting the optimal independent variables for each bus. Long et al. in Long2024 propose a clustering-based approach to select the best linear power flow model depending on the similarity of the linearization error features of different operating conditions.
The authors in ONeill2012 and Castillo2016 propose the IV-Flow model, a linear approximation of the power flow equations by formulating the active and reactive power flow equations in rectangular coordinates in terms of voltages and currents and linearizing them using a first-order Taylor series expansion. Such approximation, along with the linearization of other constraints of the AC-OPF problem, is iteratively solved until a convergence criterion is satisfied. Mhanna and Mancarella Mhanna2022 propose a sequential linear programming approach based on the construction of supporting hyperplanes and halfspaces and the exclusive solution of linear programming problems. Coffrin and Van Hentenryck Coffrin2014 propose the LPAC model, which relies on a convex approximation of the cosine terms in the power flow equations and a Taylor approximation of the remaining nonlinear terms. Bolognani and Dörfler Bolognani2015 propose an implicit linearization of the power flow manifold with a bound on the worst-case error.
On the other hand, first-order Taylor series expansions can also be applied to derive convex approximations, where all the convexities are preserved in the model while the nonconvexities are linearized. Jabr in Jabr2006 proposes a second-order conic (SOC) approximation based on his SOC relaxation, where the quadratic equality constraints are relaxed into a convex inequality and the nonconvex (arctangent) constraints are linearized. Coffrin et al. in Coffrin2012 derive a quadratic programming approximation based on a second-order Taylor series expansion of the sine and cosine functions, assuming a near-nominal voltage magnitude approximation and a small angle difference, and on the relaxation of the quadratic equality constraints into inequalities. Baradar and Hesamzadeh in Baradar2015 propose a conic approximation of the OPF problem, assuming near nominal voltage magnitudes and small angle differences, expressed in terms of voltage magnitudes, power injections, and power losses. Šepetanc and Pandžić in Sepentanc2021 propose a second-order Taylor approximation of the AC power flows in polar coordinates by distributing power losses to both branch ends, and they assess its accuracy in a unit commitment problem.
Merkli et al. Merkli2021 formulate the AC power flow problem as a nonconvex quadratically constrained linear problem aiming at minimizing the deviation of power generation from given setpoints. The problem is reformulated as a difference-of-convex functions programming problem and solved using a difference-of-convex functions algorithm. The main limitations of such work are threefold: (i) the computational complexity of splitting the indefinite matrices in the nonconvex quadratic constraints, which relies on their full eigendecomposition (which takes $\mathcal{O}(n^{3})$ time), (ii) the matrices to be split depend on the network topology, which makes such decomposition impractical in scenarios where the topology varies as it requires a new decomposition each time the topology is altered, and (iii) the convexified constraints do not preserve the sparsity of the original ones and are generally dense, which is computationally challenging for optimization solvers in terms of internal operations and memory usage.
To bridge this gap, this paper proposes a quadratically-constrained convex approximation of the AC-OPF problem (QCAC) that does not rely on typical engineering assumptions such as high reactance-to-resistance ratio, near-nominal voltage magnitudes, or small angle differences, while preserving the structural sparsity of the original nonconvex constraints. The proposed QCAC approximation is derived by lifting the AC-OPF in rectangular coordinates, which allows reformulating the problem as a difference-of-convex functions program without requiring matrix-based decompositions. The reformulated problem is then convexified around a given set of nodal voltages by linearizing only the concave components of the voltage bilinearities.
In contrast to existing difference-of-convex approaches that rely on eigendecomposition of network-dependent matrices and iterative DC algorithms Merkli2021 , the proposed method operates directly on individual bilinear voltage products, avoids dense matrix decompositions, and yields convex constraints that preserve the sparsity of the original formulation. As a result, the QCAC approximation scales favorably to large-scale networks and remains applicable under topology changes without requiring recomputation of network-dependent decompositions. The proposed approximation shows two important features: (i) it retains all the convex information of the nonconvex voltage relationships and solely linearizes their concave term, unlike other linear approximations, and (ii) it does not make any additional assumptions in terms of angle differences, reactance-to-resistance ratio, or voltage magnitudes. The accuracy and computational performance of the proposed approximation are numerically tested using realistic test systems up to 30,000 buses and two proof-of-concept case studies in transmission and distribution grids.
The remainder of this paper is organized as follows: Section 2 presents the formulation of the AC-OPF problem in rectangular coordinates and derives the proposed QCAC approximation. Section 3 illustrates the numerical performance of the proposed approximation with respect to the SOC and SDP outer approximations and a first-order Taylor series linearization. Section 4 demonstrates a case study of potential further applications of the proposed model. Section 5 concludes the paper.
2 Proposed Convex Approximation of the AC-OPF Problem
Next, we present the nomenclature and general formulation of the AC-OPF problem in lifted rectangular coordinates and derive the proposed approximation.
2.1 Nomenclature
  • Sets
  • $\mathcal{E}$
    Set of branches.
  • $\mathcal{E}_{i}$
    Set of branches connected to node $i$.
@@ 2.1 Nomenclature
  • Constants
@@
  • $\overline{N}_{i}^{\rm sh}$
    Maximum step of shunt compensation at node $i$.
  • $\overline{P}_{k}/\underline{P}_{k}$
    Maximum/Minimum active power output of unit $k$.
  • $\overline{Q}_{k}/\underline{Q}_{k}$
    Maximum/Minimum reactive power output of unit $k$.
  • $\overline{S}_{ij}$
    Maximum apparent power flow through the branch between nodes $i$ and $j$.
  • $\overline{V}_{i}/\underline{V}_{i}$
    Maximum/Minimum voltage magnitude at node $i$.
@@
  • Variables
@@
2.2 Problem Formulation
@@
2.3 The QCAC Approximation of the AC Optimal Power Flow
@@
Note that Problem (3) is an exact reformulation of the AC-OPF problem, which remains nonconvex due to the convex terms in the right-hand side of inequalities (3b)-(3f), which render these constraints nonconvex. Constraints (3b)–(3f) are written as $f^{+}(x)\leq f^{-}(x)$, where both $f^{+}$ and $f^{-}$ are convex functions. For inequalities of this form, convexity is preserved if $f^{+}(x)$ is convex and $f^{-}(x)$ is concave. Specifically, if $f^{-}(x)$ is linear or concave, the inequality defines a convex constraint, since convex functions minus concave functions lead to convex inequalities. However, if $f^{-}(x)$ is convex, the inequality will not define a convex constraint, and the problem will remain nonconvex. The proposed reformulation admits various (piecewise-)linearizations, as the nonconvexities in the right-hand side can be expressed as separable functions. A drawback of piecewise linearization is the need for SOS2 constraints, making the problem mixed-integer.
To preserve the continuous nature of the AC-OPF problem, we propose linearizing these right-hand side convex terms using a first-order Taylor series approximation around given nodal voltages, $V_{i}^{\rm re}$ and $V_{i}^{\rm im}$, effectively convexifying constraints (3b)-(3f). It is important to note that the quality of the proposed approximation heavily depends on the choice of the linearization point, as it is the only source of information used to convexify the problem. Specifically, the quality of the QCAC approximation is highly sensitive to the choice of base voltages, $V_{i}^{\rm re}$ and $V_{i}^{\rm im}$, around which the first-order Taylor series linearization is performed. For standalone applications of the QCAC model in real-time operations, a reliable linearization point can typically be obtained from the most recent solution of the state estimator. Alternatively, high-quality convexification points can be predicted using learning-based algorithms, leveraging the solutions of historical instances to predict optimal convexification points. However, the further the operating point moves from the linearization point, the larger the error in the approximation, particularly in extreme system conditions where higher-order nonlinearities become significant. To mitigate this, the choice of base voltages should be made carefully to reflect the operating conditions, ideally based on recent measurements, or predicted operating states of the system.
The proposed QCAC approximation can be modeled as follows:
@@
$\;c_{ii}\leq 2\left(V_{i}^{\rm re}v_{i}^{\rm re}+V_{i}^{\rm im}v_{i}^{\rm im}\right)-\left(\Big(V_{i}^{\rm re}\Big)^{2}+\Big(V_{i}^{\rm im}\Big)^{2}\right)+\xi_{i}^{\rm c},\;\forall i\in\mathcal{N},$ (4c)
@@
The proposed QCAC approximation can be also implemented using a sequential convex optimization procedure, summarized in Algorithm 1, which iteratively updates the convexification point and solves a sequence of convex QCQP subproblems.
Algorithm 1 Sequential QCAC approximation for AC-OPF
1:Base voltages $(V^{\rm re,0},V^{\rm im,0})$, penalty coefficient $\rho^{0}>0$, maximum penalty coefficient $\rho^{\max}$, tolerance $\varepsilon>0$, and $\mu>1$
2:Approximate AC-OPF solution $(v^{\rm re},v^{\rm im},p^{\rm g},q^{\rm g})$
3:Set $t\leftarrow 0$
4:repeat
5:  Form the QCAC subproblem (4) by linearizing around $(V^{\rm re,t},V^{\rm im,t})$ with penalty coefficient $\rho^{\rm t}$
6:  Solve the convex QCQP (4) to obtain
$\Xi^{t+1}=\big(v^{\rm re,t+1},v^{\rm im,t+1},p^{\rm g,t+1},q^{\rm g,t+1},\xi^{\rm c,t+1},\xi^{\rm c,t+1}_{ij},\xi^{\rm s,t+1}_{ij}\big)$
7:  Update base voltages:
$V^{\rm re,t+1}\leftarrow v^{\rm re,t+1},\qquad V^{\rm im,t+1}\leftarrow v^{\rm im,t+1}$
8:  Update penalty coefficient: $\rho^{\rm t+1}=\min\{\mu\rho^{\rm t},\rho^{\rm max}\}$
9:  Check stopping criterion: $\sum_{i\in\mathcal{N}}\xi_{i}^{\rm c,t+1}+\smashoperator[]{\sum_{(i,j)\in\mathcal{E}}^{}}\left(\xi_{ij}^{\rm c,t+1}+\xi_{ij}^{\rm s,t+1}\right)\leq\varepsilon$
10:  $t\leftarrow t+1$
11:until stopping criterion is satisfied
12:return $\Xi^{t}=\big(v^{\rm re,t},v^{\rm im,t},p^{\rm g,t},q^{\rm g,t},\xi^{\rm c,t},\xi^{\rm c,t}_{ij},\xi^{\rm s,t}_{ij}\big)$
2.4 Advantages of the QCAC approximation
@@
    The convexified constraints retain more information with respect to other linear approximations since they preserve all the information of the convex component of each nonconvex term and only linearize its concave part Lipp2016 . Conversely, in the linearization-based approaches, the second-order information is the objective constraints and/or constraints is lost. Figure 1 illustrates in a simple difference-of-convex (nonconvex) constraint, $y\geq x^{2}-\lvert x\rvert$, the approximate convex set obtained using the proposed convex approximation and the one obtained a standard first-order Taylor series linearization.
@@
    The QCAC approximation can be used along learning-based techniques that can leverage historical (or synthetically generated) instances to learn the mapping between nodal demands and nodal voltages, which correspond to our convexification points Pineda2024 . Notice that since the QCAC approximation is convex, it can be embedded within the training pipeline of a learning model as a differentiable layer, enabling the use of packages like DiffOpt.jl Bensancon2024 and CVXPYLayers Agrawal2019 to automate the computation of the gradients.
@@
3 Numerical experiments
In this section, we provide numerical experiments to evaluate the performance of the proposed QCAC approximation in terms of feasibility, optimality gap, and runtime. The data of the studied instances can be retrieved from the PGLib-OPF library pglib . The numerical experiments have been implemented on a Macbook Pro laptop with an Apple M2 Pro processor and 32 GB of RAM under JuMP 1.10 with Mosek 10.1.2, Ipopt, and Gurobi 11.0.2 as the optimization solvers for the convex, nonconvex, and mixed-integer nonconvex optimization problems, respectively.
3.1 Experimental setup
The results of the proposed QCAC approximation are compared with those obtained using a first-order Taylor series linearization (TS) of constraints (1l)-(1n), as well as with a second-order conic (SOC) and a semidefinite (SDP) relaxation Jabr2006 ; Bai2008 , for selected systems with up to 1354 buses. For the QCAC and TS models, the linearization/convexification points are set to the solution of the base case. For larger systems of 6,495, 13,659, and 30,000 buses, we study the performance of the QCAC model with respect to the solution of the original problem. The nonconvex, SOC, and SDP models are solved in PowerModels.jl in JuMP Coffrin2018 . For each system, we generate 100 demand samples by multiplying the base case’s active and reactive power demand by a random vector drawn from a normal distribution with mean $1$ and standard deviation $0.1$. All numerical experiments in this paper are performed on balanced network models. The proposed QCAC approximation is formulated in rectangular voltage coordinates and is not intrinsically limited to single-phase systems; extension to three-phase unbalanced distribution networks would require introducing phase-specific voltage variables and applying the same difference-of-convex decomposition and linearization procedure to the resulting phase-coupled bilinear terms.
3.2 Metrics
Optimality gap